If 5 lbs of mixed nuts containing 50% peanuts was mixed with 20 lbs of mixed nuts containing 75% of peanuts what percent of the new mixture is peanuts?
Answer by TimothyLamb(4379)

---
50*5 + 75*20 = x(5 + 20)
250 + 1500 = 25x
25x = 1750
x = 70
---
answer:
the final mixture is 70% peanuts
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php